# Generalized Nash Equilibrium Solutions in Dynamic Games With Shared Constraints

### v2

@@ 3.1 Limitations of Normalized GNE solutions
As described above, the normalized solution is usually just one of many possible solutions. In some cases, this solution may not provide a reasonable or intuitive solution. To illustrate this, we solve the problem presented in 2.1 using the normalized solution assumption. Assuming that the Lagrange multipliers of the shared constraints are identical for all players, the Lagrangian functions for the problem are:
$\begin{split}\mathcal{L}_{1}=-x_{1}+x_{2}+\frac{1}{2}v_{1}^{2}+\mu_{1}(x_{1}-x_{1}(0)-\Delta t\cdot v_{1})+\sigma(x_{2}-x_{3})\\ \mathcal{L}_{2}=-x_{2}+x_{1}+\frac{1}{2}v_{2}^{2}+\mu_{2}(x_{2}-x_{2}(0)-\Delta t\cdot v_{2})+\sigma(x_{2}-x_{3})\\ \mathcal{L}_{3}=-x_{1}+x_{2}+\frac{1}{2}v_{3}^{2}+\mu_{3}(x_{3}-x_{3}(0)-\Delta t\cdot v_{3})+\sigma(x_{2}-x_{3})\end{split}$ (12)
Applying the first-order optimality conditions - stationary:
$\begin{split}&\frac{\partial\mathcal{L}_{1}}{\partial[x_{1},v_{1}]}=\begin{bmatrix}-1+\mu_{1}\\ v_{1}-\mu_{1}\Delta t\end{bmatrix}=0\Rightarrow\mu_{1}=1\Rightarrow v_{1}=\Delta t\\ &\frac{\partial\mathcal{L}_{2}}{\partial[x_{2},v_{2}]}=\begin{bmatrix}-1+\mu_{2}+\sigma\\ v_{2}-\mu_{2}\Delta t\end{bmatrix}=0\Rightarrow\mu_{2}=1-\sigma\Rightarrow v_{2}=\Delta t(1-\sigma)\\ &\frac{\partial\mathcal{L}_{3}}{\partial[x_{3},v_{3}]}=\begin{bmatrix}\mu_{3}-\sigma\\ v_{3}-\mu_{3}\Delta t\end{bmatrix}=0\Rightarrow\mu_{3}=\sigma\Rightarrow v_{3}=\Delta t\sigma\end{split}$ (13)
By the complementary slackness condition, there are two options for a solution, either the constraint is active ($\sigma=0$), which gives the following solution:
$(x_{1},x_{2},x_{3})=(x_{1}(0)+\Delta t^{2},x_{2}(0)+\Delta t^{2},x_{3}(0))=(1,1.5,0.75)$ (14)
This solution violates the shared constraint and is therefore eliminated. The second option is that the shared constraint is active $(\sigma>0)$:
$\begin{split}&x_{2}=x_{3}\Rightarrow x_{2}(0)+\Delta t\cdot v_{2}=x_{3}(0)+\Delta t\cdot v_{3}\Rightarrow\\ &x_{2}(0)+\Delta t^{2}(1-\sigma)=x_{3}(0)+\Delta t^{2}\sigma\Rightarrow\sigma=\frac{\Delta t^{2}-x_{3}(0)+x_{2}(0)}{\Delta t^{2}}=\frac{1}{2}\Rightarrow\\ &(x_{1},x_{2},x_{3})=(x_{1}(0)+\Delta t^{2},\frac{x_{2}(0)+x_{3}(0)+\Delta t^{2}}{2},\frac{x_{2}(0)+x_{3}(0)+\Delta t^{2})}{2}=(1,1.125,1.125)\end{split}$ (15)
@@
The solution obtained from the normalized approach is counterintuitive. In the normalized solution, Car 3 does not block Car 2 as expected. Instead, it moves forward, worsening its own cost function. If Car 3 remains stationary, it would achieve a lower cost. The normalize GNE constrained by the fact that $\sigma$ is shared by all players gives a solution that car 3 would not have chosen under any other reasonable condition. It can be easily verified that any solution of the form $x_{2}=x_{3}\in[0.75,1.125]$ is a GNE. Therefore, the normalized solution is merely one possible option for Car 3, and at first glance, it is not the most appealing.
4 Non Normalized Solution For Dynamic Games
@@
The non-normalized solution of a GNEP is a solution $x\in\mathcal{X}$ such that the KKT conditions hold and the Lagrange multipliers of the shared constraints are not the same for all players. As mentioned in Section 3, the most popular tools to solve a GNEP are QVI or MCP reformulations, which produce the normalized solution when used as discussed in the previous literature. While these formulations offer many mathematical and numerical advantages, including compatibility with widely available solvers, Generalizing these formulations to allow for non-normalized solutions can be beneficial in many cases as shown by the example in section 2.1. Next, we proposes a novel method to exploit the MCP reformulation of a GNEP to calculate non-normalized solutions. Recall the KKT conditions from the original GNEP formulation (6). The main idea was to compute a single set of shared Lagrange multipliers and introduce a strictly positive scaling factor for each player’s shared constraints. Now let’s define for every player $i$ a diagonal matrix of weights:
$A_{i}\in\mathcal{D}^{+}_{m},$ (16)
$\mathcal{D}^{+}_{m}:=\{D\in\mathbb{R}^{m\times m}\mid D=$ $\text{diag}(d_{1},\cdots,d_{m}),\,d_{i}>0\,\forall i=1,\cdots,m\}.$
@@
Each $A_{i}$ is used to scale the shared Lagrange multipliers for player $i$ relative to the other players. Incorporating these factor matrices into the Lagrangina function (5) produces:
$\mathcal{L}_{i}:=J_{i}(x_{i},x_{-i})+\mu_{i}^{T}h_{i}(x_{i})+\lambda_{i}^{T}\cdot g_{i}(x_{i})+(A_{i}\sigma)^{T}\cdot s(x_{i},x_{-i})$ (17)
Thus, the KKT conditions in (6) are reformulated as:
$\nabla_{x^{i}}J_{i}(x^{i},x^{-i})+\nabla_{x^{i}}(\mu^{\top}h_{i})+\nabla_{x^{i}}(\lambda_{i}^{\top}g_{i})+\nabla_{x^{i}}((A_{i}\sigma)^{\top}s)=0,\quad\forall i\in\{1,\dots,N\},$ (18)
$0\leq\lambda_{i}\perp g_{i}\leq 0,\quad\forall i\in\{1,\dots,N\},$
@@
$\lambda_{i}\geq 0,\quad\forall i\in\{1,\dots,N\},$
$\sigma\geq 0.$
$h_{i}=0,\quad\forall i\in\{1,\dots,N\}$
@@
Given a set of matrices $\{A_{i}\}_{i=1}^{N}$ as defined in (16) and the tuple $(\bar{x},\{\bar{\mu}_{i}\}_{i=1}^{N},\{\bar{\lambda}_{i}\}_{i=1}^{N},\bar{\sigma})$ that solves the KKT conditions in (18), then $\bar{x}$ is a solution to the GNEP if a suitable constraint qualification holds.
@@
$\bar{\sigma}_{i}=A_{i}\bar{\sigma},\quad\forall i\in\{1,\dots,N\}.$ (19)
@@
The tuple $(\bar{x},\{\bar{\mu}_{i}\}_{i=1}^{N},\{\bar{\lambda}_{i}\}_{i=1}^{N},\{\bar{\sigma}_{i}\}_{i=1}^{N})$ that solves the KKT conditions in (6) and thus by Theorem 1, $\bar{x}$ is a solution of the GNEP. ∎
Theorem 2 provides a method for calculating non-normalized solutions to a GNEP using the tools developed for normalized solutions. The solution process can use the same MCP formulation as in 11 with the same available solvers (e.g., PATH), with the modified Lagrangian that includes the factor matrices. If all factor matrices $A_{i}$ are identical, the solution reduces to the normalized GNE. The factor matrices $A_{i}$ can be interpreted as representing the relative “aggressiveness” of each player: lower values in $A_{i}$ for active shared constraints indicate a reduction in the corresponding Lagrange multipliers, allowing the player to improve their cost relative to others. Notice that the diagonal entries of $A_{i}$ are strictly positive, and therefore any positive terms in $\bar{\sigma}$ (representing to active constraints) remain positive in $\bar{\sigma}_{i}$, while zero terms in $\bar{\sigma}$ remain zero in $\bar{\sigma}_{i}$. Furthermore, and without lose of generality, the factor matrix for player 1 can be defined as $A_{1}=I_{m_{0}}$. This is due to the fact that the factor matrices are all relative. The same solution $(\bar{x},\{\bar{\mu}_{i}\}_{i=1}^{N},\{\bar{\lambda}_{i}\}_{i=1}^{N},\{\bar{\sigma}_{i}\}_{i=1}^{N})$ is valid, if all factor matrices are divided by some factor and $\bar{\sigma}$ is multiplied by the same factor - this will ensure that the Lagrange multipliers $\bar{\sigma_{i}}$ will stay the same.
5 Analytical Examples
We use two examples to show the effectiveness of the method presented. The first is an example used in [7, 14]. The problem is a dynamic game between 2 player with the following setup (the variables notation are as in the original papers):
@@ 5 Analytical Examples
$\min_{x}\,(x-1)^{2}\qquad\qquad\qquad$ $\min_{y}\,\left(y-\frac{1}{2}\right)^{2}$ (20)

### v3

@@ III-A Limitations of Normalized GNE solutions
As described above, the normalized solution is usually just one of many possible solutions. In some cases, this solution may not provide a reasonable or intuitive solution. To illustrate this, we solve the problem presented in II-A using the normalized solution assumption. Assuming that the Lagrange multipliers of the shared constraints are identical for all players, the solution of the GNEP gives the following GNE:
$\begin{split}x_{1}&=x_{1}(0)+\Delta t^{2}=1.0\\ x_{2}=x_{3}=\frac{1}{2}&(x_{2}(0)+x_{3}(0)+\Delta t^{2})=1.125\end{split}$ (13)
@@
The solution obtained from the normalized approach is counterintuitive. In the normalized solution, Car 3 does not block Car 2 as expected. Instead, it moves forward, worsening its own cost function. If Car 3 had remained stationary, it would have achieved a lower cost. The normalized GNE constrained by the fact that $\sigma$ is shared by all players gives a solution the car 3 would not have chosen under any other reasonable condition. It can be easily verified that any solution of the form $x_{2}=x_{3}\in[0.75,1.125]$ is a GNE. Therefore, the normalized solution is merely one possible option for Car 3, and it is not the most appealing for team Car 1 and Car 2.
IV Non-Normalized Solution For Dynamic Games
@@
The non-normalized solution of a GNEP is a solution $x\in\mathcal{X}$ such that the KKT conditions hold and the Lagrange multipliers of the shared constraints are not the same for all players. As mentioned in Section III, the most popular tools to solve a GNEP are QVI or MCP reformulations, which produce the normalized solution when used as discussed in the previous literature. While these formulations offer many mathematical and numerical advantages, including compatibility with widely available solvers, generalizing these formulations to allow for non-normalized solutions can be beneficial in many cases as shown by the example in section II-A. Next, we propose a simple and novel method to exploit the MCP reformulation of a GNEP to calculate non-normalized solutions. Recall the KKT conditions from the original GNEP formulation (6). The main idea was to compute a single set of shared Lagrange multipliers and introduce a strictly positive scaling factor for each player’s shared constraints. Now let’s define for every player $i$ a diagonal matrix of weights:
$~~~~~~~~~~~~~~A_{i}\in\mathcal{D}^{+}_{m}$, (14)
$\mathcal{D}^{+}_{m}$$\{$ $D\in\mathbb{R}^{m\times m}\mid D=\text{diag}(d_{1},\cdots,d_{m}),$
$d_{i}\in\mathbb{R}_{++}\,\forall i=1,\cdots,m\}.$
@@
Each $A_{i}$ is used to scale the fictitious shared Lagrange multipliers for player $i$ to produce the actual Lagrange multipliers. Incorporating these factor matrices into the Lagrangian of each player $i$ gives:
$\mathcal{L}_{i}:=J_{i}(x^{i},x^{-i})+\mu_{i}^{T}h_{i}(x^{i})$ $+\lambda_{i}^{T}\cdot g_{i}(x^{i})+$ (15)
$+(A_{i}\sigma)^{T}s(x^{i},x^{-i})$ (16)
where $\sigma\in\mathbb{R}^{m_{0}}$ is the fictitious Lagrangian multipliers vector which is shared by all players.
The KKT conditions in (6) are reformulated as:
$\nabla_{x^{i}}\mathcal{L}_{i}(x^{i},x^{-i},A_{i})=0,\quad\forall i\in\{1,\dots,M\},$ (17)
$0\leq\lambda_{i}\perp g_{i}\leq 0,\quad\forall i\in\{1,\dots,M\},$
@@
$\lambda_{i}\geq 0,h_{i}=0,\quad\forall i\in\{1,\dots,M\},$
$\sigma\geq 0$
@@
Given a set of matrices $\{A_{i}\}_{i=1}^{M}$ as defined in (14) and the tuple $(\bar{x},\{\bar{\mu}_{i}\}_{i=1}^{M},\{\bar{\lambda}_{i}\}_{i=1}^{M},\bar{\sigma})$ that solves the KKT conditions in (17), then $\bar{x}$ is a solution to the GNEP if a suitable constraint qualification holds.
@@
$\bar{\sigma}_{i}=A_{i}\bar{\sigma},\quad\forall i\in\{1,\dots,M\}.$ (18)
@@
The tuple $(\bar{x},\{\bar{\mu}_{i}\}_{i=1}^{M},\{\bar{\lambda}_{i}\}_{i=1}^{M},\{\bar{\sigma}_{i}\}_{i=1}^{M})$ that solves the KKT conditions in (6) and thus by Theorem 1, $\bar{x}$ is a solution of the GNEP. ∎
Theorem 2 provides a method for calculating non-normalized solutions to a GNEP using the tools developed for normalized solutions. In order to compute the solution one can use the same MCP formulation as in (12) with the same available solvers (e.g., PATH), with the modified Lagrangian that includes the factor matrices. If all factor matrices $A_{i}$ are identical, the solution reduces to the normalized GNE. Without loss of generality, since the factor matrices $A_{i}$ do not have an absolute scale, we can impose a normalization rule to reduce the number of free parameters. For instance, the factor matrix of player 1 can be fixed as $A_{1}=I_{m_{0}}$. The same solution $(\bar{x},\{\bar{\mu}_{i}\}_{i=1}^{M},\{\bar{\lambda}_{i}\}_{i=1}^{M},\{\bar{\sigma}_{i}\}_{i=1}^{M})$ remains valid if all factor matrices $A_{i}$ are divided by a common scalar factor and $\bar{\sigma}$ is multiplied by that same factor, ensuring that the resulting Lagrange multipliers $\bar{\sigma}_{i}$ remain unchanged.
An alternative normalization rule can require that, for each shared constraint, the sum of the scaling factors across all players equals one:
$\sum_{i=1}^{N}(A_{i})_{jj}=1\qquad\forall j=1,...,m_{0}.$ (19)
Next, we provide two results that offer an interpretation of the matrices $A_{i}$ and help build intuition for the proposed formulation. The first result establishes that, under appropriate scaling (as introduced above) and the assumptions of Rosen’s setting, each solution of the GNEP corresponds to a unique set of factor matrices $A_{i}$.
Corollary 2.1.
Given a Rosen’s Setting GNEP, and if LICQ holds and given a factor matrices normalization rule then every solution of the GNEP corresponds to a unique set of factor matrices (on the set of active shared constraints) and every solution of the GNEP can be expressed by a unique set of factor matrices.
Proof of Corollary 2.1.
This corollary follows directly from Theorem 1(3) and Theorem 2. ∎
Corollary 2.1 implies that the entire solution set of the GNEP can be explored by solving for different admissible configurations of the factor matrices $A_{i}$ that satisfy the predefined normalization rule. Note, however, that not every configuration of factor matrices necessarily yields a feasible or valid equilibrium.
Building on this, the next result characterizes the effect of varying the factor matrices on the players’ costs. Specifically, Corollary 2.2 shows that, for a class of separable and convex problems with quadratic shared constraints, reducing the scaling factors of a given player relative to the others cannot worsen that player’s cost.
Corollary 2.2.
Consider Rosen’s setting of a generalized Nash equilibrium problem (GNEP) with a shared quadratic constraint. Suppose the following hold:
  1. (i)
    Each player’s cost is separable:
    $J_{i}(x_{i},x_{-i})=\phi_{i}(x_{i}),\qquad i=1,\dots,M$
    with no cross-dependence on the other player’s decision variable, and $\phi_{i}$ are strongly convex.
  2. (ii)
    The shared constraint is quadratic with symmetric block diagonal matrix $Q$,
    $s(x)=\frac{1}{2}x^{T}Qx+q^{\top}x+b=\frac{1}{2}\sum_{i=1}^{M}x_{i}^{T}Q_{ii}x_{i}+\sum_{i=1}^{M}q_{i}^{T}x_{i}+b$
  3. (iii)
    A factor-matrix normalization rule is imposed.
  4. (iv)
    $(x^{1},\dots,x^{M})$ is a solution of the GNEP corresponding to some set of factor matrices $\{A_{i}\}_{i=1}^{M}$, and The linear independence constraint qualification (LICQ) holds at the solution.
Then, reducing the values of the factor matrix $A_{i}$ of some player $i$ relative to the others will not deteriorate the cost of that player $i$.
Proof of Corollary 2.2.
For simplicity we first show the proof for corollary 2.2 for a 2 player dynamic game with a single active shared constraint:
KKT conditions of the problem and assuming that the shared constraint is active:
$F_{1}(x_{1},x_{2},\sigma,\alpha)$ $=\nabla_{x_{1}}J_{1}(x_{1},x_{2})+\alpha\sigma\,\nabla_{x_{1}}s(x_{1},x_{2})=0,$ (20)
$F_{2}(x_{1},x_{2},\sigma,\alpha)$ $=\nabla_{x_{2}}J_{2}(x_{1},x_{2})+(1-\alpha)\sigma\,\nabla_{x_{2}}s(x_{1},x_{2})=0,$
$F_{3}(x_{1},x_{2})$ $=s(x_{1},x_{2})=0.$
Differentiate $F(x,\sigma,\alpha)=0$ w.r.t. $\alpha$ and solve the bordered linear system:
$\underbrace{\begin{bmatrix}H_{11}&H_{12}&\alpha\,\nabla_{x_{1}}s\\ H_{21}&H_{22}&(1-\alpha)\,\nabla_{x_{2}}s\\ (\nabla_{x_{1}}s)^{\!\top}&(\nabla_{x_{2}}s)^{\!\top}&0\end{bmatrix}}_{\mathcal{J}}\begin{bmatrix}\dfrac{dx_{1}}{d\alpha}\\[4.0pt] \dfrac{dx_{2}}{d\alpha}\\[4.0pt] \dfrac{d\sigma}{d\alpha}\end{bmatrix}=\begin{bmatrix}-\sigma\,\nabla_{x_{1}}s\\[4.0pt] \phantom{-}\sigma\,\nabla_{x_{2}}s\\[4.0pt] 0\end{bmatrix}$ (21)
where
$H_{11}$ $=\nabla^{2}_{x_{1}x_{1}}\phi_{1}+\alpha\sigma\,Q_{11},$ $H_{12}$ $=\alpha\sigma\,Q_{12},$
$H_{21}$ $=(1-\alpha)\sigma\,Q_{21},$ $H_{22}$ $=\nabla^{2}_{x_{2}x_{2}}\phi_{2}+(1-\alpha)\sigma\,Q_{22}$
$\nabla_{x_{1}}s$ $=Q_{11}x_{1}+q_{1},$ $\nabla_{x_{2}}s$ $=Q_{22}x_{2}+q_{2}$
Define compact blocks:
$H$ $=\begin{bmatrix}H_{11}&H_{12}\\[2.0pt] H_{21}&H_{22}\end{bmatrix},\qquad g=\begin{bmatrix}\alpha\,\nabla_{x_{1}}s\\ (1-\alpha)\,\nabla_{x_{2}}s\end{bmatrix}$$\qquad$ (22)
$r$ $=\begin{bmatrix}\nabla_{x_{1}}s\\ \nabla_{x_{2}}s\end{bmatrix},\qquad\qquad b=\sigma\begin{bmatrix}\nabla_{x_{1}}s\\ -\,\nabla_{x_{2}}s\end{bmatrix}$.
Solve for $\dfrac{d\sigma}{d\alpha}$ and $\dfrac{dx}{d\alpha}=[\dfrac{dx_{1}}{d\alpha},\dfrac{dx_{2}}{d\alpha}]^{T}$ with the Schur complement. Notice that the matrix $H\succ 0$ - strong convexity of $\phi_{i}$ and the PSD of $\sigma Q_{ii}$ means that $H^{-1}$ exists and is PD:
$\frac{d\sigma}{d\alpha}=\,\frac{r^{\!\top}H^{-1}b}{\,r^{\!\top}H^{-1}g\,},\qquad\frac{dx}{d\alpha}=-\,H^{-1}\!\left(g\,\frac{d\sigma}{d\alpha}+b\right)$ (23)
Now,
$\frac{dJ_{1}}{d\alpha}=\nabla_{x_{1}}J_{1}^{\!\top}\frac{dx_{1}}{d\alpha}+$ $\nabla_{x_{2}}J_{1}^{\!\top}\frac{dx_{2}}{d\alpha}=$
$=\underbrace{\begin{bmatrix}\nabla_{x_{1}}J_{1}\quad\nabla_{x_{2}}J_{1}\end{bmatrix}}_{c^{T}}\frac{dx}{d\alpha}=c^{\!\top}\frac{dx}{d\alpha}.$
Using $F_{1}=0$ and separability, we have $\nabla_{x_{2}}J_{1}=0$ and
$\nabla_{x_{1}}J_{1}\;=\;-\,\alpha\sigma\,\nabla_{x_{1}}s.$
Hence
$c\;=\;\begin{bmatrix}\nabla_{x_{1}}J_{1}\\[2.0pt] 0\end{bmatrix}\;=\;-\,\alpha\sigma\,r_{1},\quad r_{1}:=\begin{bmatrix}\nabla_{x_{1}}s\\[2.0pt] 0\end{bmatrix},~r_{2}:=\begin{bmatrix}0\\[2.0pt] \nabla_{x_{2}}s\end{bmatrix}$.
Recall from (22) that
$r$ $=\begin{bmatrix}\nabla_{x_{1}}s\\ \nabla_{x_{2}}s\end{bmatrix}=r_{1}+r_{2},$
$g$ $=\begin{bmatrix}\alpha\,\nabla_{x_{1}}s\\ (1-\alpha)\,\nabla_{x_{2}}s\end{bmatrix}=\alpha r_{1}+(1-\alpha)r_{2},$
$b$ $=\sigma\begin{bmatrix}\nabla_{x_{1}}s\\ -\,\nabla_{x_{2}}s\end{bmatrix}=\sigma(r_{1}-r_{2}).$
Introduce the $H^{-1}$–inner product $\langle u,v\rangle=\langle u,v\rangle_{H^{-1}}:=u^{\top}H^{-1}v$ and set
$a:=\langle r_{1},r_{1}\rangle>0,\quad d:=\langle r_{2},r_{2}\rangle>0,\quad c_{0}:=\langle r_{1},r_{2}\rangle=0,$
so that $ad\geq 0$. A direct calculation gives
$r^{\top}H^{-1}b=\sigma(a-d),\qquad r^{\top}H^{-1}g=\alpha a+(1-\alpha)d.$
Therefore, from (23),
$\frac{d\sigma}{d\alpha}=-\,\sigma\,\frac{a-d}{\alpha a+(1-\alpha)d}.$
Using $\dfrac{dx}{d\alpha}=-\,H^{-1}\!\big(g\,\dfrac{d\sigma}{d\alpha}+b\big)$ and $c=-\alpha\sigma r_{1}$,
$\frac{dJ_{1}}{d\alpha}=c^{\top}\frac{dx}{d\alpha}=-\,c^{\top}H^{-1}\!\left(g\,\frac{d\sigma}{d\alpha}+b\right)=\alpha\sigma\,r_{1}^{\top}H^{-1}\!\left(g\,\frac{d\sigma}{d\alpha}+b\right).$
Substituting the identities above and simplifying yields the compact expression
Thus $J_{1}$ is non-decreasing in $\alpha$, so reducing the factor of player 1 (i.e., decreasing $\alpha$ under the normalization rule) cannot worsen player 1’s cost. This completes the proof for the two-player case with a single active shared constraint.
The same reasoning can be extended to a general $M$-player game with multiple shared constraints under the given assumptions. Specifically, we define
$r_{i}$ $=[0,\dots,\nabla_{x_{i}}s,\dots,0]\in\mathbb{R}^{n},$ (24)
$H$ $=\operatorname{blkdiag}\big(\nabla^{2}_{x_{1}x_{1}}\phi_{1}+\alpha_{1}\sigma Q_{11},\dots,$
$\qquad\qquad\qquad\dots,\nabla^{2}_{x_{M}x_{M}}\phi_{M}+\alpha_{M}\sigma Q_{MM}\big)\succ 0.$
Stacking the KKT stationarity and shared-constraint equations and differentiating with respect to $\alpha_{i}$ yields the same bordered linear system as in the two-player case. Solving it via the Schur complement gives
@@
$\frac{d\sigma}{d\alpha_{i}}=-\,\sigma\,\frac{\langle r,\,r_{i}\rangle_{H^{-1}}}{\langle r,\,g\rangle_{H^{-1}}},\qquad\frac{dx}{d\alpha_{i}}=-\,H^{-1}\!\left(g\,\frac{d\sigma}{d\alpha_{i}}+b_{i}\right),$
where $r:=\sum_{k=1}^{M}r_{k}$, $g:=\sum_{k=1}^{M}\alpha_{k}r_{k}$, and $b_{i}:=\sigma\bigl(r_{i}-\sum_{k\neq i}r_{k}\bigr)$. Because the costs are separable, $\nabla_{x}J_{i}=[0,\dots,\nabla_{x_{i}}J_{i},\dots,0]=-\alpha_{i}\sigma\,r_{i}$, so
$\frac{dJ_{i}}{d\alpha_{i}}$ $=\nabla_{x}J_{i}^{\top}\frac{dx}{d\alpha_{i}}$ (25)
$=\alpha_{i}\sigma^{2}\,\frac{\langle r_{i},r_{i}\rangle_{H^{-1}}\,\langle r,r\rangle_{H^{-1}}-\langle r_{i},r\rangle_{H^{-1}}^{2}}{\langle r,g\rangle_{H^{-1}}}\;\geq 0.$
by the Cauchy–Schwarz inequality in the $H^{-1}$ inner product, and under the block-diagonal assumption $Q_{ij}=0$ for $i\neq j$, we obtain $\langle r_{i},r_{j}\rangle_{H^{-1}}=0$ for $i\neq j$, the denominator simplifies to $\langle r,g\rangle_{H^{-1}}=\sum_{k=1}^{M}\alpha_{k}|r_{k}|_{H^{-1}}^{2}>0$, Hence, $\tfrac{dJ_{i}}{d\alpha_{i}}\geq 0$, establishing that decreasing the entries of $A_{i}$ relative to those of the other players cannot worsen player $i$’s cost in the general $M$-player case. This completes the proof. ∎
The factor matrices $A_{i}$ can be interpreted as representing the relative “aggressiveness” of each player: lower values in $A_{i}$ for active shared constraints indicate a reduction in the corresponding Lagrange multipliers, allowing the player to improve their cost relative to others. Notice that the diagonal entries of $A_{i}$ are strictly positive, and therefore any positive terms in $\bar{\sigma}$ (representing active constraints) remain positive in $\bar{\sigma}_{i}$, while zero terms in $\bar{\sigma}$ remain zero in $\bar{\sigma}_{i}$.
V Analytical Examples
We use two analytical examples to show the effectiveness of the method presented.
V-A Example 1
The first is an example used in [facchinei2007generalized, facchinei2009generalized]. The problem is a dynamic game between 2 player with the following setup (the variables notation are as in the original papers):
$\min_{x}\,(x-1)^{2}\qquad\qquad$ $\min_{y}\,\left(y-\frac{1}{2}\right)^{2}$ (26)